OpenStudy (anonymous):

In Arrhenius equation Arrhenuis constant 'A' and Ea are 4*10^-13 and 19.6*10^3 if the reaction is of first order at what temperature will the half life period will be 10 times of its initial value?

OpenStudy (frostbite):

\[\large t _{1/2}^{*}=10 ~ t _{1/2}\] \[\large \frac{ \ln(2) }{ k _{r}^{*} }=10 \frac{ \ln(2) }{ k _{r} }\] \[\large \frac{ \ln(2) }{ A ~ \exp \left( \frac{ -E _{a} }{ R ~ T _{2} } \right) }=10 ~ \frac{ \ln(2) }{ A ~ \exp \left( \frac{ -E _{a} }{ R ~ T _{1} } \right) }\] Just some notes for my self.

OpenStudy (frostbite):

Except we need information... We need to calculate the change then? solve for Delta T?

OpenStudy (frostbite):

if we had a start temperature (T1) we could solve for T2 and get an answer... but if assume that is not the case I see 2 options: 1) Substitute an expression for the temperature. 2) Solve for Delta T.

OpenStudy (anonymous):

We need atleast one temperature to solve it..

OpenStudy (anonymous):

i think we need to find answer in terms of Temperature

OpenStudy (frostbite):

All that go though my mind right now is that: \[\large \frac{ d(\ln[k _{r}]) }{ d(T ^{-1}) }=\frac{ -E _{a} }{ R }\]

OpenStudy (frostbite):

I don't got any idea how to get in our A.

OpenStudy (aaronq):

i guess you can find a temperature thats relative to the one you're at now. t(1/2)=ln2/k t(1/2)*k/ln2=constant \[\frac{ 10*t _{1/2}*k _{10} }{ \ln2 }=\frac{ t _{1/2}*k _{1} }{ \ln2 }\] \[\frac{ [10*t _{1/2}] }{ t _{1/2}}=\frac{ k _{1} }{ k _{10} } \to \frac{ k _{1} }{ k _{10} }=10\] \[ \ln( \frac{k _{1} }{ k _{10} })=\frac{ E _{a} }{ R }(\frac{ 1 }{ T _{1} }-\frac{ 1 }{ T _{10} })\] \[\ln (\frac{ k _{1} }{ k _{10} })*\frac{ R }{ E _{a} }=\frac{ 1 }{ T _{1} }-\frac{ 1 }{T _{10} }\] \[\frac{ E _{a} }{ R*\ln10 }=T _{1}-T _{10}\] this way you can get a relative temperature difference

OpenStudy (frostbite):

That or we can write out the half life as a function of temperature and determine how long it most take for the half-life to increase 10 fold: \[\large t _{1/2}(T)=\frac{ \ln(2) }{ A } \exp \left( \frac{ -E _{a} }{ RT } \right)\] A good old exponential function.

OpenStudy (aaronq):

sounds good !

OpenStudy (aaronq):

i'm just wondering, when you subbed in k into the first order equation, did you make sure e^(-Ea/RT) was at the bottom? \[t _{1/2}=\frac{ \ln2 }{ A*e ^{\frac{ -Ea }{ RT }} }\]

OpenStudy (frostbite):

Hmmm woops, then rewrite to perhaps? \[\large t _{1/2}(T)=\frac{ \log(2) }{ A } ~ \exp \left( \frac{ E _{a} }{ RT } \right)\]

OpenStudy (aaronq):

thats works haha